Circular No. 8851

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V390 NORMAE D. K. Lynch, R. W. Russell, R. J. Rudy, and T. Prater, The Aerospace

Corporation; and C. E. Woodward, University of Minnesota, report 0.8to 5.5-μm spectroscopy of V390 Nor (cf. IAUC 8850) using the Infrared Telescope Facility’s 3-m telescope (+ SpeX) on June 21 UT. V390 Nor is confirmed as a classical nova well into its emission-line phase. While the spectrum is dominated by low-excitation lines of H I, C I, N I, O I, and Ca II, He I features are already evident. Fe II emission is present but is not strong. Weak P-Cyg profiles extending to 1800 km/s from line center are present on the H I, He I, and O I lines and on the very strong Ca II infrared triplet. The reddening derived from the strong O I lines is E(B − V) ≈ 1.0, and there is no evidence of dust in the nova shell.

SUPERNOVAE 2007cl, 2007cm, 2007cn, 2007co, 2007cp, 2007cq Six additional apparent supernovae have been found on unfiltered CCD

frames and given designations: 2007cl by T. Puckett and G. Sostero (cf. IAUC 8843); 2007cm by W. Kloehr (cf. IAUC 8789); 2007cn and 2007cp by W. Li and H. Khandrika (LOSS/KAIT; cf. IAUC 8843); 2007co by J. Nicolas (Vallauris, France); and 2007cq by T. Orff and J. Newton (cf. IAUC 8794). Discovery observations:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007cl May 23.34 17^{h}48^{m}21.^{s}19 +54^{o}09′05′′.2 17.6 3′′.2 W, 8′′.2 N
2007cm May 24.02 12 42 45.18 +55 08 57.1 16.1 21′′.5 E, 13′′.5 N
2007cn May 29.48 22 13 55.79 +13 45 23.9 18.0 2′′.4 W, 3′′.9 N
2007co June 4.98 18 23 03.60 +29 53 49.6 16.6 8′′ E, 15′′ S
2007cp June 13.20 12 42 12.72 −17 24 07.7 17.7 3′′.8 E, 5′′.5 N
2007cq June 21.34 22 14 40.43 + 5 04 48.9 15.8 3′′.3 W, 5′′.7 N

Additional information is available on Central Bureau Electronic Telegrams: 2007cl in NGC 6479, CBETs 972, 974 (type-Ic, discovered a few days past maximum); 2007cm in NGC 4644, CBETs 973, 974 (young, type-IIn supernova); 2007cn in UGC 11953, CBET 975; 2007co in MCG +05-43-16, CBETs 977, 978 (type-Ia, discovered well before maximum); 2007cp in IC 807, CBETs 980, 981 (type-Ia); 2007cq, CBET 983. Note that 2007ck (cf. IAUC 8843) was discovered in MCG +05-43-16 only a couple of weeks before 2007co was found.

2007 June 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green